OpenStudy (anonymous):

volume of sphere is 288pi cm^3 Diameter=

OpenStudy (anonymous):

The volume is given by (4/3)pi r^3. So: 288pi=(4/3)pi r^3. Solve for r. Giving: r=6 The diameter is twice the radius so d=12.

OpenStudy (anonymous):

i have to solve 288pi=(4/3)pi r^3 first

OpenStudy (anonymous):

Divide by pi. 288=(4/3) r^3 Divide by (4/3) (same as multiplying by 3/4 288(3/4)=(3/4)(4/3) r^3 216=r^3 r=cuberoot(216) r=6.